PRINTER VERSION BACK NEXT Assume that Current Designs will have sales of $3.60 million with two-thirds of the sales dollars in rotomoulded kayaks and one-third of the sales dollars in composite $805,000 of fixed costs are allocated to the rotomoulded kayaks and $196,040 to the composite kayaks, prepare a CVP income statement for each product line. Rotomoulded Kayaks Composite Kayaks LINK TO TEXT LINK TO TEXT LINK TO TEXT LINK TO TEXT LINK TO TEXT LINK TO TEXT Using the information in part (d), calculate the degree of operating leverage for each product line. (Round answers to 2 decimal places,e.g.52.75.) Rotomoulded Kayaks Composite Kayaks Degree of operating leverage LINK TO TEXT LIK TO TEKT LINK TO TEXT LINK TO TEXT LINK TO TEXT LUNK TO TEXT
Question7 Current Designs manufactures two different types of kayaks: rotomoulded kayaks and composite kayaks. The following information is available for each product line. Rotomoulded Composite 1,200 $2,000 7201,340 Sales price/unit Variable costs/unit The companys fixed costs are $1,001,040. An analysis of the sales mix identifies that rotomolded kayaks make up 80% of the total units sold. Determine the weighted-average unit contribution margin for Current Designs. Weighted-average unit contribution margins 516 LINK TO TEXT LINK TO TEXT uNK TO TEXT LINK TO TEXT LINK TO TEXT LINK TO TEXT Determine the break-even point in units for Current Designs and identify decimal places, e.g.5,275.) how many units of each type of kayak will be sold at the break-even point. (Round answers to o Break-even sales 1940 units Rotomoulded Kayaks Composite Kayaks 413 units 103 units
Answer #1

Q. 7 CVP Income Statement

Rotomoulded Kayakas Composite Kayakas
Sales $       2,400,000 $       1,200,000
Variable Costs $       1,440,000 $          804,000
Contribution Margin $          960,000 $          396,000
Fixed Costs $          805,000 $          196,040
Net Income $          155,000 $          199,960

No. of Units sold of Rotomoulded Kayakas = $2,400,000 / 1,200 = 2000 units & of Composite Kayakas = $1,200,000 / 2,000 = 600 units

Variable cost of Rotomoulded Kayakas - 2,000 * $720 = $1,440,000 & of Composite Kayakas - 600 * $1,340 = $804,000

Degree of Operating Leverage = Contribution Margin / Net Income

Degree of Operating Leverage of Rotomoulded Kayakas = 960,000 / 155,000 = 6.19

Degree of Operating Leverage of Composite Kayakas = 396,000 / 199,960 = 1.98.